1 (i) Determine the plinth area for the building forming with center line dimensions of outer walls 8.6m X 14.6m and the thickness of the wall 0.40m.

A. 132.61m2
B. 135.57m2

C. 113.96m2
D. 135m2

(ii) The routing coefficients for the Muskingum method to route the flood in a given channel are asC, = —0.35,C„ = 0.45. Then C,would be
(a) 0.55

(b) -0,8

(c) 0.65

d) 0.9

(iii) Froude number is used-
(a) To determine the frictional value for object moved from air.

(b) To determine the viscosity for the object

(c) To determine the resistance for an object moved from water.

(d) To determine the lower value of the fluid to resist by the load itself.
Answer #1

1) ii)

In muskingham method,

Sum of the coefficients is equal to 1 so

C0 + C1 + C2 = 1

Two of them are given so other can be found out

C2 = 1 - (C1 +C0​​​​​​​​)

= 1 - 0.45 + 0.35

= 0.9 i.e. option (d)

1) iii)

Froud number is used to determine the resistance offered for an object moved from water..

i.e. answer is option (C)